### v1

@@ 3 Results and Analyses
082.5+11.3 | NGC 6833 | 297.444 | 48.961 | 2086764737566091520 | T | 0.3 | 1.0 | – | 978 d |
061.4-09.5 | NGC 6905 | 305.596 | 20.104 | 1816547660416811008 | T | 21.65 | 1.0 | – | 700 d? |
084.6-07.9 | Kn 26 | 320.789 | 38.97 | 1965268354805776896 | T | 55.0 | 1.0 | – | 230 d? 1.18 hr |
044.1+05.8 | CTSS 2 | 282.695 | 12.625 | 4504643171308707072 | T | – | 0.99 | – | – |
034.3+06.2 | K 3-5 | 277.941 | 4.086 | 4283434037501296128 | T | 5.0 | 0.99 | – | – |
228.2-22.1 | LoTr 1 | 88.778 | -22.901 | 2917223705359238144 | T | 71.0 | 0.78 | 11.28 yr (1), 6.48 d (2) | 6.48 d |
069.6+15.7 | Kn J1857.7+3931 | 284.426 | 39.517 | 2103243049609815552 | P | 75.0 | 1.0 | – | – |
032.5-03.2 | K 3-20 | 285.542 | -1.813 | 4262011019866089984 | T | – | 0.98 | – | – |
014.0-02.5 | Tan 2 | 276.321 | -17.941 | 4096278711670069248 | T | 3.0 | 0.83 | – | – |
045.0-12.4 | WHTZ 3 | 299.497 | 4.792 | 4248349003725092864 | T | 46.0 | 0.99 | – | – |
044.3-04.1 | IPHAS J192717.94+081429.4 | 291.825 | 8.241 | 4296182638654047232 | P | – | 1.0 | – | – |
@@ 3.2.3 Discussion
The light curves also bear a resemblance to another class of phenomena: Long Secondary Period (LSP) photometric modulations observed in several RGB and post-AGB stars (for example, see Kiss et al. 2006, Nicholls et al. 2009, and Soszyński et al. 2021). The phenomenon is not yet well understood and several possibilities have been proposed over the past couple of decades. One is related to binarity, where the companion is a low-mass object (large planet or brown dwarf), which carries a dust cloud around it, causing the occultation of the star (Soszyński et al., 2021). In such a scenario, we would expect an infrared excess in the spectral energy distribution (SED) of the CSPN. Interestingly, strong infrared excess is seen in both NGC 6905 and Kn 26 (Figure 7). However, this excess is unlikely to be caused by a localized dust cloud around a companion, as the dimensions of the cloud required to match the observed excess are large. For example, the mid- and far-infrared excess in NHC 6905 is matched with a blackbody of temperature of K and radius of AU. This still does not rule out this possibility as the dust cloud causing the photometric variation might just be a smaller embedded component.
@@ 3.3.2 Follow-up Optical Spectroscopy
The LRS2 instrument is described in detail by Chonis et al. (2014, 2016). Briefly, LRS2 provides integral-field-unit (IFU) spectroscopy with 280 -diameter lenslets that cover a field of view (FOV) on the sky. LRS2 is composed of two arms: blue (LRS2-B) and red (LRS2-R). All of our observations are made with the target placed in the LRS2-B FOV. The LRS2-B arm employs a dichroic beamsplitter to send light simultaneously into two spectrograph units: the “UV” channel (covering 3640–4645 Å with a resolving power of 1910), and the “Orange” channel (covering 4635–6950 Å with a resolving power of 1140). Data reduction was performed using the pipelines developed by G.R.Z.888https://github.com/grzeimann/Panacea,999https://github.com/grzeimann/LRS2Multi (Zeimann et al., in prep.).
@@ 3.3.5 Spectral Energy Distribution: Infrared Excess
To look for the presence of dust in these systems, we investigate their spectral energy distributions (SED). We retrieved the SED of CTSS 2 and K 3-5 (around ′′ from the target coordinate) from the VizierR online photometry viewer tool. 101010http://vizier.cds.unistra.fr/vizier/sed/ The result is presented in Figure 13. Some scatter is evident in the data, especially for K 3-5, which is possibly due to the extended nature of the objects leading to suboptimal photometry. Nevertheless, the presence of significant infrared excess in both the objects is evident.

### v2

@@ 3 Results and Analyses
| 082.5+11.3 | NGC 6833 | 297.4440 | 48.9611 | 2086764737566091136 | T | 0.3 | 1.0 | – | 978 d |
| 061.4-09.5 | NGC 6905 | 305.5958 | 20.1045 | 1816547660416810880 | T | 21.65 | 1.0 | – | 700 d? |
| 084.6-07.9 | Kn 26 | 320.7889 | 38.9700 | 1965268354805776896 | T | 55.0 | 1.0 | – | 230 d? 1.18 hr |
| 044.1+05.8 | CTSS 2 | 282.6953 | 12.6249 | 4504643171308707072 | T | – | 0.99 | – | – |
| 034.3+06.2 | K 3-5 | 277.9409 | 4.0859 | 4283434037501295744 | T | 5.0 | 0.99 | – | – |
| 228.2-22.1 | LoTr 1 | 88.7775 | -22.9007 | 2917223705359238016 | T | 71.0 | 0.78 | 11.28 yr (1), 6.48 d (2) | 6.48 d |
| 069.6+15.7 | Kn J1857.7+3931 | 284.4260 | 39.5167 | 2103243049609815296 | P | 75.0 | 1.0 | – | – |
| 032.5-03.2 | K 3-20 | 285.5423 | -1.8126 | 4262011019866089856 | T | – | 0.98 | – | – |
| 014.0-02.5 | Tan 2 | 276.3213 | -17.9413 | 4096278711670069248 | T | 3.0 | 0.83 | – | – |
| 045.0-12.4 | WHTZ 3 | 299.4972 | 4.7921 | 4248349003725092608 | T | 46.0 | 0.99 | – | – |
| 044.3-04.1 | IPHAS J192717.94+081429.4 | 291.8248 | 8.2415 | 4296182638654046592 | P | – | 1.0 | – | – |
@@ 3.2.3 Discussion
The light curves also bear a resemblance to another class of phenomena: Long Secondary Period (LSP) photometric modulations observed in several RGB and post-AGB stars (for example, see Kiss et al. 2006, Nicholls et al. 2009, and Soszyński et al. 2021). The phenomenon is not yet well understood and several possibilities have been proposed over the past couple of decades. One is related to binarity, where the companion is a low-mass object (large planet or brown dwarf), which carries a dust cloud around it, causing the occultation of the star (Soszyński et al., 2021). In such a scenario, we would expect an infrared excess in the spectral energy distribution (SED) of the CSPN.888We retrieved the SED of CTSS 2 and K 3-5 (around ′′ from the target coordinate) from the VizierR online photometry viewer tool http://vizier.cds.unistra.fr/vizier/sed/. We rejected all data points for which no errors are quoted. The SEDs for CTSS 2 and K 3-5 (discussed in Section 3.3) have been retrieved in the same way. Interestingly, strong infrared excess is seen in both NGC 6905 and Kn 26 (Figure 7). However, this excess is unlikely to be caused by a localized dust cloud around a companion, as the dimensions of the cloud required to match the observed excess are large. For example, the mid- and far-infrared excess in NHC 6905 is matched with a blackbody of temperature of K and radius of AU. This still does not rule out this possibility as the dust cloud causing the photometric variation might just be a smaller embedded component.
@@ 3.3.2 Follow-up Optical Spectroscopy
The LRS2 instrument is described in detail by Chonis et al. (2014, 2016). Briefly, LRS2 provides integral-field-unit (IFU) spectroscopy with 280 -diameter lenslets that cover a field of view (FOV) on the sky. LRS2 is composed of two arms: blue (LRS2-B) and red (LRS2-R). All of our observations are made with the target placed in the LRS2-B FOV. The LRS2-B arm employs a dichroic beamsplitter to send light simultaneously into two spectrograph units: the “UV” channel (covering 3640–4645 Å with a resolving power of 1910), and the “Orange” channel (covering 4635–6950 Å with a resolving power of 1140). Data reduction was performed using the pipelines developed by G.R.Z.999https://github.com/grzeimann/Panacea,101010https://github.com/grzeimann/LRS2Multi (Zeimann et al., in prep.).
@@ 3.3.5 Spectral Energy Distribution: Infrared Excess
To look for the presence of dust in these systems, we investigate their spectral energy distributions (SED). The result is presented in Figure 13. Some scatter is evident in the data, especially for K 3-5, which is possibly due to the extended nature of the objects leading to suboptimal photometry. Nevertheless, the presence of significant infrared excess in both the objects is evident.